OpenStudy (anonymous):

Write in point-slope form the equation of the line that passes through the given point and has the given slope (-8, -2 m= 2

OpenStudy (anonymous):

y=2x+18

OpenStudy (anonymous):

y+2=2(x+8) y=2x+18